OpenStudy (anonymous):

SOLVE BY QUADRATIC FORMULA: X^2-5X=7

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

?? wher is second equetion?

OpenStudy (mohsen):

\[-b \pm \sqrt{b ^{2}-4ac}/2a\] a=1, b=-5, c=-7

OpenStudy (mohsen):

\[-5\pm \sqrt{(-5)^{2}-4(1)(-7)}/2(1)\] \[-5\pm \sqrt{25+28}/2\] \[-5\pm \sqrt{53}/2\] \[i) -5+\sqrt{53}/2\] = find yourself! \[ii) -5-\sqrt{53}/2\] = find yourself!